OpenStudy (anonymous):

Joe has a collection of nickels and dimes that is worth $6.20. If the number of dimes was doubled and the number of nickels was decreased by 7, the value of the coins would be $10.35. How many nickels and dimes does he have?

OpenStudy (anonymous):

$6.20 = 620 cents a: number of nickels b: number of dimes \[5a+10b = 620\] \[5(a-7)+10(2b)=1035\]\[5a-35+20b=1035\]\[5a+20b=1070\] Two equations \[5a+20b=1070\]\[5a+10b=620\] Subtract. \[10b=450\]\[b=45\] Plug in b to get a. \[5a+450=620\]\[5a=170\]\[a=34\] a=34, b=45